OpenStudy (anonymous):

Identify the axis of symmetry and vertex of f(x) = –2x^2 –2x–1. Axis of symmetry: x = 0.5; Vertex: (0.5, – 0.50) Axis of symmetry: x = – 0.5; Vertex: (– 0.5, 0.50) Axis of symmetry: x = – 0.5; Vertex: (– 0.5, – 0.50) Axis of symmetry: x = 0.5; Vertex: (– 0.5, – 0.50)

OpenStudy (anonymous):

@fashionismybesty

OpenStudy (anonymous):

http://www.mathway.com/graph/NTUxODk this is an amazing site for graphing

OpenStudy (freckles):

Do you know how to write in vertex form? If involves completing the square? \[f(x)=-2x^2-2x-1 \\ f(x)=-2(x^2+x)-1 \\ f(x)=-2(x^2+x+L)-1+2L \\ \text{ notice } -2L+2L=0 \\ \text{ we need to figure out what } L \text{ needs to be to } \\ \text{ complete the square inside that one ( )}\]

OpenStudy (triciaal):